You will probably be pleased to hear that MEG is now fully operational
again.
We have also taken time to get the system recalibrated, eye trackers
installed and a complete overhaul of the mechanics of the bed.
The noise level is lower as the source of the small electrical pulse
artefacts has been identified and removed.
All MEG channels are working.
The original problem was due to a faulty interface between the fibre
optics taking data out of the shielded room into the data acquisition
rack. The board has been replaced.
Bookings can now be taken for MEG.
You may also be pleased to hear that MRI has had a routine maintenance
and all is well.
happy scanning!
--
Gary Green
York Neuroimaging Centre
The Biocentre
York Science Park
Innovation Way
Heslington
York
YO10 5DG
http://www.ynic.york.ac.uk
tel. 01904 435349
fax 01904 435356
mobile 07986 778954